# Information That Matters: Exploring Information Needs of People Affected by Algorithmic Decisions

### v1

@@ 3. Method
To elicit information needs of affected stakeholders, we conducted an interview study in German language with 24 participants recruited from a local job agency and a local apartment complex inhabited mainly by retirees (Section 3.4). We presented participants with one of two ADM use cases (Section 3.1): an employment prediction algorithm or a health wristband for elderly care. Participants were tasked to vote whether the presented use cases should be adopted in general. Before deciding, participants were asked to gather information on the systems by asking questions in two 15-minute inquiry phases and received verbal explanations by the main author of this paper in response. The study closed with an interview at around 50 minutes. Interviews were analyzed using a combination of inductive and deductive analysis (as described in Section 3.3).
With this procedure, we aimed to fulfill multiple study goals: firstly, we introduced a task that motivated inquiry of information (i.e., voting on adoption), secondly, by first letting participants inquire freely and flexibly about the use cases and second providing the XAI Question Bank (Liao et al., 2020) as reference we could examine both intuitive information needs as well as the bank’s utility for their inquiry, and third, by specifically eliciting participants’ understanding and perceptions we gave room for articulations about aspects not covered by the inquiry itself. The study setup is described in Section 3.2 and the procedure is depicted in Figure 1.
@@ 3.1. Use cases
Choice of use cases. Use cases were selected to adhere to the definition of ”high-risk” ADM systems (Commission, 2021) while differing in several dimensions, including: a) domain of use (employment and health), b) degree of tangibility (the wristband is a physical device, while the AMS algorithm does not have a physical counterpart), c) user groups, and d) functionality. These differences were intended to allow for examination of their influence on participants’ information needs (RQ1), understanding (RQ2), and perceptions of the systems (RQ3). As the two use cases are representative of two types of well-known ADM systems that have already been implemented in comparable forms (Niklas et al., 2015; Scott et al., 2022; Kutsarova, 2020; Sahlen et al., 2018), our study aims to provide insight on how to approach explanation design for the future deployment of similar systems.
@@
Use Case A: AMS algorithm. The AMS algorithm 333AMS stands for the Public Employment Agency (Arbeitsmarktservice). is a scoring system meant to predict the employability of job-seekers in Austria. Developed between 2015 and 2021 by a private company for the Austrian Public Employment Agency and piloted in 2018, it was never used as a live system and was put on hold in 2021 due to privacy objections (Allhutter et al., 2020b). The system was meant to use job-seekers’ personal features such as age, education, and nationality444The full list of features is given in the supplementary material. to compare them with the time it took people with similar features to find employment in the past. This comparison would produce a short-term and long-term employment score for each job-seeker, which in turn would be shown to the employment agency’s personnel consultant as a decision aid for suitable support measures. Personnel consultants would be able to override the system’s scores, but would need to give a reason for doing so (Allhutter et al., 2020b; Holl et al., 2018).
The AMS algorithm has been academically and publicly discussed with respect to the risks and benefits of its deployment, including detailed reports about the system’s potential to reproduce intersectional discrimination (Allhutter et al., 2020b; Allhutter, 2021; Scott et al., 2022; Lopez, 2019). As the implementation of such a system could have large effects on job-seekers and as comparable systems have been developed in other countries (Niklas et al., 2015; Scott et al., 2022), the AMS algorithm presented a suitable use case for our study.
Use Case B: Health wristband for elderly care. The wristband is an adaptation to wearable alarm systems already in use in care organizations (Chang et al., 2023; Kutsarova, 2020; Kaun and Taranu, 2020; Sahlen et al., 2018). While existent alarm systems often require the wearer to push a button to alert care personnel, the health wristband in our study was described to continuously evaluate movement and acceleration data and raise an alarm when it detected that the wearer tripped or fell. The system would thereby not require manual activation by the wearer, making it viable for use in private homes. Further, the wristband collects data on the wearer’s sleeping, movement, and eating behavior, which the care facility could monitor and use to improve the organization of care.
3.2. Study setup
3.2.1. Scenario and decision task
@@
Participants were initially informed that they would vote over the ADM system’s adoption after the inquiry phases, introducing an incentive to gain information. After participants were given a consent form and a questionnaire on demographical information and previous knowledge, participants were presented with the decision scenario, which included a brief description of the use case and its planned deployment. The description rendered the participants as part of a small citizen referendum asked to provide their vote on the system. A small box was printed below the scenario description for participants to fill in ”yes” or ”no” after the inquiry phases, acting like a voting ballot. The design was inspired by ballots used in popular votes in Switzerland (Council, 2023).555Popular votes in Switzerland require the voter to write ”Yes” or ”No” on the ballot, which can be argued to be more deliberate compared to a tick box.
@@ 3.2.2. Inquiry phases I & II
After being briefed about the decision scenario, participants had 30 minutes in total to ask questions to the study examiner and receive verbal explanations about the ADM use case at hand before taking the decision. The time in which participants could inquire about the use cases was split into two phases of 15 minutes each: In the first phase, participants inquired about the use case intuitively, having received no initial information beyond the scenario description. In the second phase, participants were provided with a reduced version of the XAI Question Bank (Liao et al., 2020)666The amount of questions compiled in the XAI Question Bank (Liao et al., 2020) would have made its application in the study impractical due to the time necessary to scan and understand all items. Unsuitable questions were thus omitted and similar questions summarized. A detailed tabular explanation for the reduction is given in the supplementary material., a collection of questions on AI systems sourced from AI developers and designers, which they could use as a reference for further questions. As the original XAI Question Bank includes a comprehensive collection of questions, we reduced it by summarizing questions that aim towards similar topics. During the inquiry, participants were provided with a pen and paper to take notes.
3.2.3. Verbal explanations
Participants received verbal explanations from the study examiner as responses to their questions. The explanations were based on information prepared in advance and available to the study examiner during the study. For each use case, two pilot studies were conducted to collect a stock of questions that would likely be asked in the subsequent studies and which served as guidance for preparing information. When questions could not be answered due to a lack of information, they were re-formulated to inquire about how and why the requested information would be relevant to the information needs of participants (for examples of interview excerpts, please refer to the supplementary material).
@@ 3.2.4. Self-reported understanding, confidence, and risks & benefits
Participants reported their understanding, decision confidence, and perceptions of the system’s risks and benefits (and an explanation for their answer) at two points: first, before the inquiry phases, and second, after making the decision. Self-reported understanding and decision confidence were elicited through 5-point Likert scales. To encourage participants to speak about the trade-offs between the systems’ risks and benefits, these were split into two dimensions: societal and personal. For each elicitation of risks and benefits, two visual analogue scales (10 cm) (Rugg and Petre, 2007) were combined to form a two-dimensional Cartesian coordinate system with the origin located at 5 cm on both axes. By denoting perceived risk on the x-axis and perceived benefit on the y-axis, the coordinate system formed four quadrants of varying risk and benefit combinations (as depicted in Figure 3). Participants drew marks on the axes that were measured and then rounded to the first decimal place.
3.3. Analysis
@@
For RQ1-Information Needs, we use thematic analysis (Braun and Clarke, 2006) to examine participants’ information needs both inductively and deductively using the XAI Question Bank (Liao et al., 2020). All questions raised by participants in the two inquiry phases were coded according to the category of inquiry. Categories already present in the XAI Question Bank (Liao et al., 2020) were assigned deductively, and topics not yet covered were defined inductively from the data. The XAI Question Bank (Liao et al., 2020) thus served as a starting point for a comparative analysis between participants’ information needs and those of AI practitioners, highlighting similarities and discrepancies between the groups. An intermediary version of the coding was checked by three coders for inter-coder agreement by independent analysis of the transcripts and a subsequent comparison of code categories.
For RQ2-Understanding, we evaluated participants’ responses on the self-reported understanding and decision confidence Likert scales, and used a deductive analysis of explanatory stances and understanding challenges as outlined in Section 2. For RQ3-Perception, we analyze how explanations provided during the inquiry phases impact participants’ perceptions of the systems’ risks and benefits, and how these perceptions in turn influence which questions are asked.
@@ 3.4. Participants
Table 1 provides an overview of the study participants. We recruited participants from two main locations:In the employment prediction use case, we collaborated with the local employment agency Job-TransFair, which established contact with both job-seekers and personnel consultants. In the wristband use case, we used street sampling in an apartment complex known for the quality of living it offers to retirees.777While we contacted many institutions active in geriatric care work, none wanted to collaborate. The reasons for this are unknown. However, the difficulty of recruiting retirees for study participation over official channels is noteworthy. The recruitment criteria were i) no or little previous experience with either algorithms or AI systems (reported in the questionnaire) and ii) identification with the ”affected stakeholder” group as described below. All interviews were conducted in person either in office spaces or cafés, except for one conducted online. Participants were compensated with 12€ in cash, which was disclosed only after completion of the study so as not to incentivize participation solely for compensation.
All participants of our study can be defined as affected stakeholders (Langer et al., 2021) to varying degrees. In the employment prediction use case, job-seekers were defined as affected if they were searching for employment at the time of the interview or had been searching in the past. Personnel consultants fulfilled multiple stakeholder roles: first, all personnel consultants also had been job-seekers in the past, and second, being the system’s end users, its deployment would affect their interpersonal relations and collaboration with job-seekers. In the wristband use case, the same categories applied regarding geriatric care. Here, no participants were in care currently or in the past, but all had close relatives in care. Further, the age of participants in this use case was required to be around 60 years or older (B4 and B9 are the exception, as their domain expertise as nurses proved relevant to compare information needs and understanding).
| ID | Age | Education | Occupation | Decision | ID | Age | Education | Occupation | Decision |
| A1 | 54 | University | Personnel consultant | Yes | B1 | 62 | A level | Retired | Yes |
| A2 | 61 | Vocational university | Personnel consultant | No | B2 | 68 | A level | Retired | Yes |
| A3 | 54 | Secondary school | Job-seeking | Yes | B3 | 57 | Secondary school | Retired | Yes |
| A4 | 43 | A level | Personnel consultant | Yes | B4 | 46 | Apprenticeship | Geriatric care | No |
| A5 | 27 | Apprenticeship | Job-seeking | No | B5 | 80 | Secondary school | Retired | Yes |
| A6 | 53 | University | Personnel consultant | No | B6 | 68 | University | Retired | Yes |
| A7 | 63 | University | Personnel consultant | n/a | B7 | 58 | University | Freelancer | Yes |
| A8 | 56 | Apprenticeship | Job-seeking | No | B8 | 70 | Vocational university | Retired | Yes |
| A9 | 50 | A level | Job-seeking | No | B9 | 36 | University | Geriatric care | Yes |
| A10 | 63 | Secondary school | Cleaning service | No | B10 | 66 | Secondary school | Retired | Yes |
| A11 | 40 | A level | Personnel consultant | Yes | B11 | 89 | Apprenticeship | Retired | Yes |
@@ 4. Results
In this section, we present our results, structured according to our research questions: information needs of affected stakeholders (RQ1-Information Needs, Section 4.1), participant understanding and decision confidence (RQ2-Understanding, Section 4.2), and perceived risks and benefits (RQ3-Perceptions, Section 4.3).
| Category | Subcategory | Question | Information provided | ||
| Data |
Type |
What kind of data was the system trained on?* |
Content and structure of training data |
Source |
What is the source of the training data?* |
Collection and storage of training data |
Attributes |
Are the data correct / representative / safe? |
Quality, biases, safety of training data |
| System context |
Intention |
What is the intention of deploying it? |
@@
Design idea and underlying motivation |
How does the deployment process work? |
Process of preparing the system for operation |
Consequences |
What are the consequences after deployment? |
Possible effects of deployment |
Development |
How is the system developed? |
Development process and institution |
Who is the system’s intended target group? |
Scope and features of the target group |
Responsibility |
Who is responsible for the system’s deployment? |
Responsible actors |
Perspective |
What do other people think about the system? |
Votes and considerations of others |
Values |
How are moral values considered in the system? |
Alignment with moral principles |
| System usage |
Tool |
How is the system operated? |
Usage in practice, everyday application |
@@
How will the system impact interpersonal relations? |
Interpersonal work, substitution of people |
@@
How is the system integrated into existing structures? |
Interfaces, processes, organization’s influence |
Misuse |
How can the system be misused? |
Scenarios of misuse |
Example |
@@
How would the system handle [this case]? |
Specific example walkthrough |
Costs |
What are the costs? |
Financing, personal costs |
| System specs. |
Features |
@@
What features does the system consider and why?* |
Information fed into system |
Functionality |
@@
What is the scope of the system’s capability?* |
Supported range of functions |
Output |
@@
What kind of output does the system give?* |
Information received from system |
Learning |
How does the system learn?* |
Automated improvement of system |
Logic |
What is the system’s overall logic?* |
General process overview |
Algorithm |
What kind of algorithm was used?* |
Specific model used |
Mistakes |
@@
What kind of mistakes is the system likely to make?* |
Expected errors |
Limitations |
What are the limitations of the system?* |
Known boundaries |
Reliability |
How reliable are the predictions?* |
Dependability of output |
Reflexive |
@@
What does [a machine learning concept] mean? |
ML concept information |
4.1. RQ1-Information Needs: What are affected stakeholders’ information needs when deciding over the adoption of an ADM system?
@@
To address RQ1, we coded participants’ questions raised in the inquiry phases according to reoccurring categories. An overview and diagram are provided in Figure 2, a tabular representation in Table 2. The categorization covers the categories data, system context, system usage, and system specifications, each containing subcategories with a summarizing question. The subsequent sections (4.2, 4.3) describe how this information influenced participants’ decisions on the system’s adoption. The bracketed numbers denote the total number of questions asked in each subcategory.

### v6

@@ 3. Method
To elicit the information needs of AI novices, we conducted an interview study in German with 24 participants recruited from a local job agency and a local apartment complex inhabited mainly by retirees (Section 3.4). We used a qualitative approach for this study to allow for comprehensive conversations with participants and the collection of data that could be analyzed thematically (Braun and Clarke, 2006). The interviews were thus designed to develop a theoretical account of information needs that is grounded in empirical observation (Urquhart et al., 2009) (Section 3.3). The focus of this study is therefore on the conceptualization and exploration of AI novices’ information needs, which aims to lay the groundwork for further qualitative and quantitative examination in future work.
We presented participants with one of two algorithmic decision-making (ADM) use cases (Section 3.1): an employment prediction algorithm or a health wristband for geriatric care. Participants were tasked to decide whether the presented use cases should be adopted. Before deciding, participants had the opportunity to gather information about the systems by asking questions in two 15-minute inquiry phases. In response to their questions, they received verbal explanations from the main author of this paper. The study closed with an interview at around 50 minutes. The data was analyzed using inductive and deductive analysis (as described in Section 3.3).
The procedure was designed to fulfill multiple study goals. Firstly, the decision task was meant to give participants a sense of importance for their decision and to motivate them to acquire relevant information. Secondly, by letting participants inquire freely and flexibly about the use cases in the first phase and providing the XAI Question Bank (Liao et al., 2020) only in the second phase, we could examine both participants’ intuitive information needs and the bank’s impact on their inquiry. Third, by eliciting participants’ understanding of the ADM systems and their perceptions of the risks and benefits, we can examine the interactions between these aspects and participants’ information needs. The whole study setup is described in Section 3.2, and the study procedure is depicted in Figure 1.
@@ 3.1. Use cases
The two use cases were selected as representative forms of so-called “high-risk” ADM systems, i.e., systems that pose risks to the health, safety, or fundamental rights of people (European Commission, 2021). In particular, they are examples of the domains employment and biometrics, which are defined as areas with high-risk potential in the upcoming EU AI Act (European Commission, 2021).
Both use cases are thus examples of more general application areas of high-risk algorithmic systems. While models, data, and context are specific to each use case, the overarching themes of information inquiry and the utility of question-driven explanations are broadly applicable. They can be transferred to the explanation design for other algorithmic systems. In the following, each use case is briefly described, and its selection is motivated.
3.1.1. Use Case A: AMS (employment prediction) algorithm
@@
Description. The AMS algorithm 444AMS stands for the Public Employment Agency (Arbeitsmarktservice). is a scoring system meant to predict the employability of job-seekers in Austria. It was developed by a private company for the Austrian Public Employment Agency between 2015 and 2021 but was never used as a live system and was put on hold in 2021 due to privacy objections (Allhutter et al., 2020b). To predict employability chances, the system used a logistic regression model trained on historical data to calculate how job-seekers’ demographic features influenced their chances (such as age, education, nationality, etc.555The full list of features is given in the supplementary material.). This model would produce a short-term and long-term employment score for each job-seeker. The scores would be given to the job-seeker’s counselor at the employment agency to aid in deciding about suitable support measures. Counselors could overwrite the system’s group assignment of job-seekers but would need to give a reason for doing so (Allhutter et al., 2020b; Holl et al., 2018).
Choice of use case. Algorithmic tools for the profiling of job-seekers have been deployed in various countries, including Germany (Bundesagentur für Arbeit, 2021), Austria (Allhutter et al., 2020b), Poland (Niklas et al., 2015), and the Netherlands (Desiere and Struyven, 2021). The AMS algorithm is thus an example of a more general trend: the introduction of automation and algorithmic procedures into public employment agencies to increase productivity and efficiency (Allhutter et al., 2020a). Profiling tools like the AMS algorithm assist in assessing job-seekers and resource allocation and often split the responsibility for tasks between the algorithmic system and caseworkers (Scott et al., 2022). However, the introduction of these applications also repeatedly highlights which parts of human interaction cannot be easily replaced by automation, such as “the desire to be seen as a whole human being” (Scott et al., 2022). As the AMS algorithm, therefore, conjoins typical procedures, interactions, and challenges of ADM systems in public employment, it presented a suitable use case for our study.
3.1.2. Use Case B: Health wristband for geriatric care
Description. The health wristband is an assistance tool for the long-term care of older adults, which tracks and evaluates movement and acceleration data and raises an alarm when the wearer trips or falls. The system does not require manual activation by the wearer but relies on trained machine learning models to detect events (erratic movement, heart rate anomalies, etc.) in the recorded data, making it viable for use in private homes. Further, the wristband collects data on the wearer’s sleeping, movement, and eating behavior, allowing for remote monitoring of patients. The wristband thus fulfills a two-part purpose for the care organization: First, harmful events can be detected immediately, and thus, little time is lost when dispatching emergency services. Second, patients’ data can be used for the organization of care services by making overall coordination more efficient. Systems with comparable functionality have reportedly been employed in Sweden since 2016 (Kaun and Taranu, 2020; Sahlen et al., 2018).
Choice of use case. Wearable devices for identifying and monitoring patients are already in use in many care organizations (Kutsarova, 2020; Kaun and Taranu, 2020; Sahlen et al., 2018), such as nursing homes (Chang et al., 2023) and psychiatric wards (Nahavandi et al., 2022). Wristbands, in particular, can capture signals such as heart rate and body temperature while tracking movement and sleep (Seneviratne et al., 2017). This data can then be used to detect anomalies in breathing, physical movement, or other vital signs and trigger the appropriate reaction. This facilitates health status monitoring without keeping patients in the hospital or nursing homes, making them useful for various scenarios (Kutsarova, 2020). Besides being an example of wearable devices already well-established in care contexts, the wristband also incorporates a direct trade-off between perceived privacy and safety, as it can help detect risky situations in exchange for continuous monitoring (Kutsarova, 2020). Consequently, some users might reject these health devices due to feeling under surveillance (Nahavandi et al., 2022). A further known issue is prediction inaccuracy for certain population groups due to biases in the training data (Gerke et al., 2020). As the wristband is thus representative of known benefits and risks of wearable devices in care contexts, we selected it as the second use case for our study.
3.2. Study procedure and setup
The study procedure consisted of several steps that are depicted in Figure 1. This section describes the use cases and decision task, the inquiry phases I and II, the question-driven explanations during the study and the reporting mechanisms for understanding, confidence, and risks and benefits.
3.2.1. Use case and decision task
@@
At the beginning of the study, participants were informed that they would have to vote over the adoption of the ADM system after the inquiry phases, which introduced an incentive to gain information. After participants were given a consent form and a questionnaire on demographical information and previous knowledge, they were presented with the decision use case, including a brief description of the use case and its planned deployment. The description rendered the participants as part of a small citizen referendum asked to provide their vote on the system. A small box was printed below the use case description for participants to fill in “yes” or “no” after the inquiry phases, acting like a voting ballot. The design was inspired by ballots used in popular votes in Switzerland (Swiss Federal Council, 2023).666Popular votes in Switzerland require the voter to write “Yes” or “No” on the ballot, which can be argued to be more deliberate compared to a tick box.
@@ 3.2.2. Inquiry phases I & II
After being briefed about the decision use case, participants had 30 minutes to ask the study examiner questions and receive verbal explanations about the ADM use case before making their decision. The time available to participants for inquiries about the use cases was split into two phases of 15 minutes. In the first phase, participants asked questions about the use case from their own intuition. In the second phase, participants were provided with a reduced version of the XAI Question Bank (Liao et al., 2020)777A detailed tabular explanation for the reduction is given in the supplementary material., a collection of questions on AI systems sourced from AI developers and designers, which they could use as a reference for further questions. As the original XAI Question Bank includes a comprehensive collection of questions, we reduced it by summarizing questions that aim toward similar topics. Participants were given a pen and paper to take notes during their inquiry.
3.2.3. Question-driven verbal explanations
Participants received verbal explanations from the study examiner as responses to their questions. The explanations were based on information prepared in advance and available to the study examiner during the study (as depicted in Figure 2). For each use case, two pilot studies were conducted to collect a stock of questions that would likely be asked in the subsequent studies and which served as guidance for preparing information. When questions could not be answered due to a lack of information, the examiner asked how and why the requested information would be relevant to participants (for examples, please refer to the supplementary material).
@@ 3.2.4. Self-reported understanding, confidence, and risks & benefits
Participants reported their understanding, decision confidence, and perceptions of the system’s risks and benefits (and an explanation for their answer) at two points: first, before the inquiry phases, and second, after making the decision. Self-reported understanding and decision confidence were elicited through 5-point Likert scales. To encourage participants to speak about the trade-offs between the systems’ risks and benefits, these were split into societal and personal dimensions. For each elicitation of risks and benefits, two visual analog scales (10 cm) (Rugg and Petre, 2007) were combined to form a two-dimensional Cartesian coordinate system with the origin located at 5 cm on both axes. By denoting perceived risk on the x-axis and perceived benefit on the y-axis, the coordinate system formed four quadrants of varying risk and benefits combinations (as depicted in Figure 4). Participants drew marks on the axes, measured, and then rounded to the first decimal place.
3.3. Interview method and analysis
Interview method. The study employed two interview techniques throughout the study procedure (Figure 1). In inquiry phases I and II, participants were free to ask any question about the ADM use case that they found relevant. This led to a form of unstructured interview, in which not the examiner but the participant was driving information acquisition. At this point, the examiner only asked questions to clarify participants’ inquiries and to determine why they asked for certain information. In the last part of the study, the examiner conducted a semi-structured interview to gather participants’ responses on three key aspects: their overall reasoning and strategy for asking questions, their perception of the XAI Question Bank’s (Liao et al., 2020) usefulness, and their perception of whether they acquired a complete picture of the system. At this point, the study examiner followed up on thematic aspects that arose during the interview and further connected participant responses to the quantitative items elicited before to facilitate qualitative exploration (Weiss, 1995).
@@ 3.3. Interview method and analysis
Analysis. Regarding RQ1-Information Needs, we use thematic analysis (Braun and Clarke, 2006) to examine participants’ information needs inductively and deductively using the XAI Question Bank (Liao et al., 2020). All questions raised by participants in the two inquiry phases were coded according to the category of inquiry. Categories already present in the XAI Question Bank (Liao et al., 2020) were assigned deductively, and topics not yet covered were defined inductively from the data. The XAI Question Bank (Liao et al., 2020) thus served as a starting point for a comparative analysis between participants’ information needs and those of AI practitioners, highlighting similarities and discrepancies between the groups. A team of three coders from the authors’ research group checked an intermediary version of the coding for inter-coder agreement by independently analyzing the transcripts and comparing code categories.
For RQ2-Understanding, we evaluated participants’ responses on the self-reported understanding and decision confidence Likert scales and combined this with a deductive analysis using the frameworks of “explanatory stances”(Keil, 2006) and understanding challenges (Wiggins and McTighe, 2005; Keil, 2006) (described in Section 2).
For RQ3-Perception, we analyze how explanations provided during the inquiry phases impact participants’ perceptions of the systems’ risks and benefits by comparing their reports on the risks and benefits coordinate system (Section 3.2.4) before and after inquiry and analyzing their corresponding articulations. Further, we analyze how participants’ perceptions influence which questions are asked by comparing their reports on risks and benefits and information needs between use cases.
@@ 3.4. Participants
Recruitment. Table 1 provides an overview of the study participants. We recruited participants from two main locations: In the employment prediction use case, we collaborated with the local employment agency Job-TransFair, which established contact with job-seekers and personnel counselors. In the wristband use case, we used street sampling in an apartment complex known for the quality of living it offers to retirees.888While we contacted many institutions active in geriatric care work, none wanted to collaborate. The reasons for this are unknown. However, the difficulty of recruiting retirees for study participation over official channels is noteworthy. The recruitment criteria were i) no or little previous experience with either algorithms or AI systems (reported in the questionnaire) and ii) identification with either the domain expert or decision subject stakeholder group as described below. All interviews were conducted in person, either in office spaces or cafés, except for one conducted online. Participants were compensated with € 12 in cash, which was disclosed only after completion of the study so as not to incentivize participation solely for compensation.
Choice of participants. In both use cases, participants were selected such that they were representative for one of two roles: domain experts or decision subjects. Following the XAI stakeholder framework (Barredo Arrieta et al., 2020; Langer et al., 2021), we define domain experts as people who are competent in the field that the ADM system is used in (e.g., employment and healthcare) and who would be its likely users. As domain experts, we recruited personnel counselors for the employment prediction use case, and care workers and a doctor for the health wristband use case. All domain experts had multiple years of experience in their field and could readily imagine how the corresponding ADM system would integrate with their work.
We describe decision subjects as people who are impacted by an automated decision and might seek ways to contest or change it but who are not necessarily users of the system (Alfrink et al., 2023; Yurrita et al., 2022). We recruited job-seekers and people who had previously been job-seeking as decision subjects for the employment prediction use case, as well as retired people and their relatives for the health wristband use case. Most of these participants already had experience with systems comparable to our use cases, such as system providing statistical competence analyses of job-seekers and wearable alarm systems for retired people in geriatric care, which, however, did not include machine learning.
| ID | Age | Education | Occupation | Decision | ID | Age | Education | Occupation | Decision | |
| A1 | 54 | University | Personnel counselor | Yes | B1 | 62 | A level | Retired | Yes | |
| A2 | 61 | Vocational university | Personnel counselor | No | B2 | 68 | A level | Retired | Yes | |
| A3 | 54 | Secondary school | Job-seeking | Yes | B3 | 57 | Secondary school | Retired | Yes | |
| A4 | 43 | A level | Personnel counselor | Yes | B4 | 46 | Apprenticeship | Geriatric care | No | |
| A5 | 27 | Apprenticeship | Job-seeking | No | B5 | 80 | Secondary school | Retired | Yes | |
| A6 | 53 | University | Personnel counselor | No | B6 | 68 | University | Retired | Yes | |
| A7 | 63 | University | Personnel counselor | n/a | B7 | 58 | University | Freelancer | Yes | |
| A8 | 56 | Apprenticeship | Job-seeking | No | B8 | 70 | Vocational university | Retired | Yes | |
| A9 | 50 | A level | Job-seeking | No | B9 | 36 | University | Geriatric care | Yes | |
| A10 | 63 | Secondary school | Cleaning service | No | B10 | 66 | Secondary school | Retired | Yes | |
| A11 | 40 | A level | Personnel counselor | Yes | B11 | 89 | Apprenticeship | Retired | Yes | |
@@ 4. Results
In this section, we present our results, structured according to our research questions: information needs of AI novices (RQ1, Section 4.1), participant understanding and decision confidence (RQ2, Section 4.2), and perceived risks and benefits (RQ3, Section 4.3).
| Category | Subcategory | Question | Information provided |
@@
| System context | Intention | What is the intention of deploying it? | Design idea and underlying motivation |
| Deployment process | How does the deployment process work? | Process of preparing the system for operation | |
| Consequences | What are the consequences after deployment? | Possible effects of deployment | |
| Development | How is the system developed? | Development process and institution | |
| Target group | Who is the system’s intended target group? | Scope and features of the target group | |
| Responsibility | Who is responsible for the system’s deployment? | Responsible actors | |
| Ethical considerations | Is the system’s deployment right? | None, answer relies on personal assessment | |
| System usage | Tool | How is the system operated? | Usage in practice, everyday application |
@@
| Sociotechnical component | How will the system impact interpersonal relations? | Interpersonal work, substitution of people | |
@@
| Part of an organization | How is the system integrated into existing structures? | Interfaces, processes, organization’s influence | |
| Misuse | How can the system be misused? | Scenarios of misuse | |
@@
| Example | How would the system handle [this case]? | Specific example walkthrough | |
| Costs | What are the costs? | Financing, personal costs | |
| Data | Type | What kind of data was the system trained on?* | Content and structure of training data |
| Source | What is the source of the training data?* | Collection and storage of training data | |
| Attributes | Are the data correct / representative / safe? | Quality, biases, safety of training data | |
@@
| System specs. | Features | What features does the system consider and why?* | Information fed into system |
@@
| Functionality | What is the scope of the system’s capability?* | Supported range of functions | |
@@
| Output | What kind of output does the system give?* | Information received from system | |
| Learning | How does the system learn?* | Automated improvement of system | |
| Logic | What is the system’s overall logic?* | General process overview | |
| Algorithm | What kind of algorithm was used?* | Specific model used | |
@@
| Errors | What kind of mistakes is the system likely to make?* | Expected errors | |
| Limitations | What are the limitations of the system?* | Known boundaries | |
| Reliability | How reliable are the predictions?* | Dependability of output | |
@@
| Reflexive | What does [a machine learning concept] mean? | ML concept information |
4.1. RQ1-Information Needs: What information do AI novices who could be affected by algorithmic decisions need to decide about adopting an ADM system?
@@
To address RQ1, we coded participants’ questions raised in the inquiry phases to identify reoccurring question categories. The found categories are system context, system usage, data, and system specifications, and are each composed of several prototypical questions. An overview is provided in Figure 3, a tabular representation in Table 2. In the following, each of the four question categories is described and split into subcategories. We denote the total number of questions asked in each subcategory in brackets.
4.1.1. System context: Intention, deployment process, consequences, development, target group, responsibility, and ethical considerations.
This category captures questions that elicit background information on the system’s development and deployment (69 questions total).
Intention: What is the intention of deploying it? Questions about the intention of the system’s deployment were asked for both use cases (18 total), with a majority in employment prediction. These questions aimed at the underlying reasoning for introducing the system. Here we differentiate i) practical intention, such as data collection, “How did they come up with the idea of GPS? What’s the intention behind it?” (B4); and ii) structural intention, such as societal benefit, “What is the benefit for society? What did they expect from this algorithm?” (A1).
Deployment process: How does the deployment process work? Questions about the general deployment process were primarily asked in the employment prediction use case (13 of 19 total). This includes inquiries about how the system would be tested and how comparable systems were deployed in other countries.
Consequences: What are the consequences after deployment? Questions on what would follow from the system’s deployment were asked in low counts in both use cases (9 total). Topics of inquiry included the effect on the available workforce and care ratio and tended to focus on risks and challenges:
Won’t I lose labor potential if I only decide based on educational background and so on? If this algorithm is introduced, will it not lead to the loss of workforce? (A1)
Development: How is the system developed? Questions on the system’s technical development emerged mainly in the employment prediction use case (8 of 9 total). They covered, e.g., the contracted company, duration of development, and the collaboration between developers and domain experts. Personnel counselors primarily asked these questions, who related the information to their experience with contracted software developers and compared it to known procedures.
Target group: Who is the system’s intended target group? Questions on the system’s target groups emerged mainly in the wristband use case (8 of 10 total) and covered, e.g., the target group’s physical mobility and financial situation. In contrast, the two inquiries in the employment prediction use case considered target groups more on a societal level: “Would the system be applied for every population group? Can we say that every citizen would then be covered?” (A11).
Responsibility: Who is responsible for the system’s deployment? These questions (4 total) addressed responsibility for the system’s deployment, including the political decision, financing, project management, and technical implementation.
Ethical considerations: Is the system’s deployment right? Questions in this subcategory addressed perceived social wrongs or difficult trade-offs on personal and societal levels, which might be aggravated by the ADM system’s introduction (9 total). Personal topics included perceptions of unfairly distributed resources, perceived inaccessibility to own data, and over-reliance on care personnel for patients’ safety (“If I know that the person falls all the time, what do I do? If I let them walk around anyway, am I then responsible?” (B1)). Societal topics included the (in)sensibleness of sanctioning job-seekers, reflections on whether an ADM system could be more profitable if it did not aim for people’s well-being and the (in)validity of quantifying people’s attributes (“Do you believe that you can categorize every human being in such a scientific way? By some gradations and variables?” (A4)).
This category captures participants’ information needs about how the system would be used in practice and as a concrete application. Questions on usage were asked frequently in both use cases (89 questions in total by over 20 participants).
Usage as a tool: How is the system operated? Questions on practical usage details were almost exclusively gathered in the wristband use case (30 of 31 total). These typically concern how the wristband would be used, including waterproofing, charging, operation, etc. Questions like these would likely be covered by a manual accompanying the wristband.
Usage as a sociotechnical component: How will the system impact interpersonal relations? Questions in this category asked who would use the system and how the system would influence relations between people. These were gathered mostly in the employment prediction use case (13 of 14 total). Participants’ inquiries focused on how the system would affect interaction between counselor and job-seeker and covered both the counselor’s and job-seeker’s personal experience as well as their relationship:
Usage as part of an organization: How is the system integrated into existing structures? Questions that focused on the system’s interaction with existing organizational structures were present in both use cases (28 total) and covered a wide range of topics, including data processing, data access, information flow, and contestability: “What are the options for correction? Is there a system loop that handles this?” (A7).
Examples: How would the system handle [this case]? Questions in this subcategory aimed to examine how the system would behave given a specific example case and emerged in low count in both use cases (6 total). Specifically, several participants were interested in how the algorithm would handle their case.
Costs: What are the costs? Questions about costs pertained to both personal costs (“What would it cost me?” (B5)) as well as organizational expenses (“If the state pays, what is the cost? Is the money well invested?” (A4)). This subcategory appeared more often in the wristband use case (6 of 8 total).